OpenStudy (anonymous):

A certain radioactive isotope has a half-life of 5 days. If one is to make a table showing the half-life decay of a sample of this isotope from 32 grams to 1 gram; list the time (in days, starting with t = 0) in the first column and the mass remaining (in grams) in the second column, which type of sequence is used in the first column and which type of sequence is used in the second column?

OpenStudy (campbell_st):

ok.. this is exponential growth and decay is the model you are using similar to \[A = A_{0}e^{-kt}\] A0 is the initial quantity t = time and k = decay constant...?

OpenStudy (campbell_st):

is that the model..?

OpenStudy (anonymous):

i guess i honestly dont know

OpenStudy (anonymous):

i know it has something to do with arthmetic

OpenStudy (campbell_st):

well the solution will require logarithms... do you understand them..?

OpenStudy (anonymous):

ya sort of and this question is under the topic of Geometric Sequences

OpenStudy (anonymous):

The nth term of a geometric sequence has the form: an = a1r n-1 Where r is the common ratio of consecutive terms of the sequence.

OpenStudy (campbell_st):

ok... so you know a = 1 and \[a_{5} = 0.5\] and t = 5 so you need to find r the common ratio.. \[0.5 = 1 \times r^{5 -1}\] so \[r = \sqrt[5]{0.5}\] I'd expect you will need to evaluate r. then you can find t \[1 = 32r^{t - 1}\] which will require logs to solve for t